### v2

# ResGS: Residual Densification of 3D Gaussian for Efficient Detail Recovery

Abstract
Recently, 3D Gaussian Splatting (3D-GS) has prevailed in novel view synthesis, achieving high fidelity and efficiency. However, it often struggles to capture rich details and complete geometry.
Our analysis reveals that the 3D-GS densification operation lacks adaptiveness and faces a dilemma between geometry coverage and detail recovery. To address this, we introduce a novel densification operation, residual split, which adds a downscaled Gaussian as a residual. Our approach is capable of adaptively retrieving details and complementing missing geometry. To further support this method, we propose a pipeline named ResGS. Specifically, we integrate a Gaussian image pyramid for progressive supervision and implement a selection scheme that prioritizes the densification of coarse Gaussians over time. Extensive experiments demonstrate that our method achieves SOTA rendering quality. Consistent performance improvements can be achieved by applying our residual split on various 3D-GS variants, underscoring its versatility and potential for broader application in 3D-GS-based applications.
Figure 1: We propose ResGS, a pipeline to boost the rendering quality of 3D-GS [24] while improving efficiency. (a) The 3D-GS densification operations, split and clone, rely on binary selection and cannot adaptively address the challenges they encounter. (b) Our ResGS employs a single densification operation, residual split, that adaptively addresses under-reconstruction and over-reconstruction. (c) 3D-GS tends to generate Gaussians with sizes in small variation and low rendering quality. (d) From the same initialization, our method can capture fine details and retrieve sufficient geometry effectively while reducing redundancy.
1 Introduction
Figure 2: An illustration of the limitation of split and clone. (a) selects between split and clone based on a high threshold, (b) a low threshold, and in (c) the densification operation is replaced as residual split. We provide visualizations of the final rendered image. A high threshold ensures sufficient structural coverage but causes over-reconstruction, leading to blurry rendering. Conversely, a low threshold mitigates blurry rendering but results in under-reconstruction, compromising coverage and creating a difficult trade-off. Our method, however, adaptively tackles both issues without facing this trade-off.
Novel View Synthesis (NVS) has continued to be a hot topic throughout the years. Traditional methods like point clouds and meshes [4, 14, 9] can achieve high rendering speed but often lack fidelity. On the other hand, neural volumetric approaches [31, 1, 32] can achieve high fidelity for novel views but significantly sacrifice rendering efficiency. Recently, 3D-GS [24] has shown both high fidelity and fast rendering speed for novel view synthesis. However, 3D-GS still suffers from setbacks, often failing to capture fine details and recover complete geometry, as shown by red and orange rectangles in Fig. 1 and Fig. 2.
While various approaches [48, 12, 50, 43, 13] have been proposed to tackle these challenges and have significantly advanced the rendering quality of 3D-GS-based methods to SOTA levels, they have not modified the densification operation of 3D-GS. In this paper, we take a novel approach by revisiting and refining the densification operation to improve reconstruction quality.
The densification operation used in 3D-GS-based methods consists of two operations: split, designed to address over-reconstruction (where Gaussians cover too much area, resulting in a lack of detail), and clone, aimed at under-reconstruction (missing geometry). However, selecting between the two operations is challenging. 3D-GS-based methods rely on a threshold-based decision mechanism. As illustrated in Fig. 1 (a), Gaussians whose scale $s=max(s_{x},s_{y},s_{z})$ is larger than a predefined threshold $\tau_{s}$ undergo split, while those below it use clone. Due to its simplicity, the selected operation might be suboptimal and bring a dilemma in threshold selection. As highlighted by the red rectangles in Fig. 2 (a), a high threshold results in insufficient split, which prevents the model from capturing fine details, leading to blurry artifacts. On the other hand, a low threshold reduces clone operations, leading to missing geometry, as illustrated in the orange boxes in Fig. 2 (b). This dilemma limits the rendering quality. Furthermore, 3D-GS tends to first split and then clone, resulting in objects being fitted with Gaussians of similar scales. As highlighted by the blue rectangles in Fig. 1, 3D-GS fits textureless areas with an excessive number of small Gaussians, which is redundant.
To tackle the aforementioned problems, we propose residual split, a novel densification operation. As illustrated in Fig. 1 (b), the core idea is that for any under-optimized Gaussian, we spawn a downscaled replica as a residual component while reducing the original Gaussian’s opacity. This simultaneously expands spatial coverage and injects finer details, eliminating the split/clone dilemma.
However, directly applying residual split does not produce significant improvements. This is due to the inefficiency of initialization in 3D-GS. At early iterations, the scene is represented by a few coarse Gaussians that face both over-reconstruction and under-reconstruction. Gaussians added through densification are challenging to optimize, as they must address both issues simultaneously. It’s desired to develop an approach that decouples coverage and detail optimization.
Correspondingly, we propose a coarse-to-fine training pipeline, ResGS, that progressively shifts the focus of training, facilitating optimization. Firstly, we structure the training process into stages, employing multi-scale images for gradual supervision, ensuring that the early stages focus on the overall structure and the later stages focus on details. Secondly, we develop a progressive Gaussian selection scheme for densification, which gradually encourages coarser Gaussians to undergo further refinement. Therefore, finer structures are introduced at later iterations, avoiding premature optimization. These two techniques simplify optimization and thereby improve rendering quality.
Our contributions are summarized as follows:
  • We propose a novel densification operation for 3D Gaussians based on a residual strategy to overcome the limitations of fixed threshold for determining split or clone in 3D-GS;
  • We develop a coarse-to-fine training pipeline, ResGS, equipped with residual split, which incorporates an image-pyramid-based supervision schedule and a Gaussian selection scheme to enhance rendering quality;
  • Our method achieves SOTA rendering quality in NVS on three datasets, including Mip-NeRF360 [1], Tanks&Temples [25], and DeepBlending [18]. Furthermore, applying residual split to multiple 3D-GS variants consistently brings performance improvements, demonstrating that our new densification operation is more effective than densification with a fixed scale threshold.
2 Related Work
2.1 Novel View Synthesis
Novel view synthesis (NVS) aims to acquire novel view images for a scene through a set of multi-view images. As a pioneer, NeRF [31] introduced an MLP-based implicit representation. More specifically, it uses an MLP to predict the color and opacity of a point in a scene and uses differentiable volume rendering to render images. Due to its photo-realistic quality, it has been extensively studied [7, 22, 6, 47, 27, 8, 1, 2]. However, NeRF-based methods often suffer from long training and rendering time. Therefore, many works [33, 34, 19, 20] have been proposed to extend NeRF into real-time rendering.
Recently, 3D-GS [24] has stirred an evolution in the field by achieving both high-fidelity and real-time rendering. Due to its high performance, many works have enabled 3D-GS into many downstream tasks such as SLAM [30, 41], 3D generation [10, 37, 44, 52], 4D scene modeling [42, 39, 16, 26], and surface reconstruction [21, 15].
However, the original 3D-GS still has limitations, and myriad works have been proposed to increase the render quality and efficiency of 3D-GS. Mip-Splatting [46] introduced a 2D dilation filter and a 3D smoothing filter to mitigate aliasing and artifacts caused by changes in focal length. Scaffold-GS [29] proposed the concept of neural Gaussians which is to tie Gaussians around an anchor and use an MLP to predict its attributes. Thus, the model is more robust to significant view changes and can improve rendering quality. AbsGS [43] discovered that when accumulating the view space gradients used in densification through pixels, there could be a collision, for gradients at various pixels might have different positivities. They applied an absolute function before summing through pixels to mitigate the issue. Pixel-GS [50] analyzed 3D-GS artifacts from the pixel perspective and introduced a method to select Gaussians for densification based on this analysis. GaussianPro [11] introduced a progressive propagation strategy to propagate Gaussians to under-initialized areas. To reduce redundancy and increase rendering quality, Mini-Splatting [13] proposed a densification and simplification algorithm to reorganize Gaussians’ spatial positions.
With the emergence of numerous works focused on optimizing NeRF and 3D-GS, progressive training schedules—proven effective in multiple scenarios [3, 17, 36, 23] have earned attention.
2.2 Progressive Training in NVS
Several progressive training schedules have been explored upon NeRF-based methods. BungeeNeRF [40] captured different levels of details in large city-scale datasets by gradually adding closer views to the training views and appending residual blocks to the MLP throughout the training process. Pyramid NeRF [51] proposed a progressive subdivision of the scene while utilizing an image pyramid to gradually add high-frequency information to the scene. BARF [28] proposed a scheduling approach to progressively incorporate higher frequency components in positional encoding, enabling the training of NeRFs without requiring camera poses.
Recently, with the prevalence of 3D-GS, many progressive training pipelines have been proposed. Octree-GS [35] proposed an Octree-based 3D-GS approach, and the Octree depth is progressively deepened. FreGS [48] designed a progressive frequency regulation for 3D-GS which is to calculate the loss of rendered images in the frequency domain and gradually add higher frequency components.
3 Our Method
Figure 3: Overview of our ResGS pipeline. (a) The core of our pipeline, residual split, involves adding a downscaled replicate and then reducing the opacity of the original Gaussian. (b) We assign initial Gaussians a temporary attribute $l_{i}=0$ for densification selection, which is discarded after training. Next, the pipeline is split into $L$ ($L=3$) stages, with each stage trained on images downscaled using an image pyramid. Furthermore, we divide a single stage evenly into $K$ ($K=3$) substages for use in selecting Gaussians to densify. (c) The points selected for densification are determined by the substage $k$, $l_{i}$ and viewspace gradients of Gaussians. residual split is performed upon selected Gaussians, and $l_{j}$ would be assigned to new Gaussians.
3.1 Preliminaries
3D Gaussian Splatting.
The 3D-GS work[24] represents a 3D scene with a set of anisotropic 3D Gaussians for fast and high-quality rendering. Each Gaussian primitive consists of two parameters, mean $\boldsymbol{\mu}$ and covariance matrix $\mathbf{\Sigma}$, representing its position and shape:
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})}.$ (1)
To model the positive semi-definite property of the covariance matrix $\mathbf{\Sigma}$, a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ is used as:
$\mathbf{\Sigma}=\mathbf{RS}\mathbf{S}^{T}\mathbf{R}^{T},$ (2)
where a 3D vector $\mathbf{s}$ is used for the scaling matrix $\mathbf{S}$.
Furthermore, each Gaussian also contains a spherical harmonics coefficients $SH$ and opacity $o$. In rendering, tile-based rasterization is applied: 3D Gaussians $G(\mathbf{x})$ are projected as 2D Gaussians $G^{\prime}(\mathbf{x})$ on the image plane. A view-dependent color $c$ is then derived from $SH$, and the 2D Gaussians are sorted by depth and rendered using alpha-blending:
$C(\mathbf{x})=\sum_{i\in N_{\mathbf{x}}}c_{i}\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$ (3)
$\alpha_{i}(\mathbf{x})=c_{i}G_{i}^{\prime}(\mathbf{x})$
where $\mathbf{x}$ is a pixel position in the rendered image, $C(\mathbf{x})$ the color of the pixel, and $N_{\mathbf{x}}$ the number of Gaussians that covers $\mathbf{x}$.
During the training process, 3D Gaussians are periodically densified. 3D-GS calculates the average viewspace gradients $\Delta L_{i}$ over a certain number (e.g. 100) of iterations to determine whether to perform densification. If $\Delta L_{i}$ exceeds threshold $\tau$, a split or clone action is triggered. The split operation addresses over-reconstruction, where excessive space is covered, while clone addresses under-reconstruction, where additional spatial coverage is needed. To determine whether to split or clone, 3DGS[24] sets a threshold $\tau_{s}=kR$, where $k$ is a predefined factor and $R$ is the scene extent. Split occurs when the scale $s=max(\mathbf{s}_{x},\mathbf{s}_{y},\mathbf{s}_{y})$ of a Gaussian surpasses a threshold $\tau_{s}$. Clone is applied when $s<\tau_{s}$.
3.2 ResGS
We propose ResGS, a coarse-to-fine pipeline to boost the rendering quality of 3D-GS with a novel residual densification operation. Fig. 3 shows the overview of our method. We propose a novel densification operation, residual split, (Sec. 3.2.1), to tackle the limitations of the 3D-GS densification operation. Furthermore, we introduce a split-stage supervision schedule to add information to the scene gradually (Sec. 3.2.2). We also design a Gaussian selection scheme that progressively encourages densification for coarse Gaussians to enhance details, as described in Sec. 3.2.3.
3.2.1 Residual Split
The core idea of our densification operation is to generate downscaled Gaussians as residuals to supplement those requiring densification. The newly added Gaussian adaptively enhances the reconstructed model, allowing the scene to be refined as needed. The method is named as residual split. As illustrated in Fig. 3 (a), for a to-be-densified Gaussian $G_{i}$, our densification operation consists of two steps. Firstly, we add a down-scaled replicate $G_{j}$. Specifically, its opacity $o_{j}$, rotation $\mathbf{R}_{j}$, spherical harmonics coefficients $SH_{j}$ would be same as $G_{i}$, $\boldsymbol{\mu}_{j}$ is sampled using the original 3D Gaussian as a PDF for sampling, and $\mathbf{S}_{j}$ is defined as $\mathbf{S}_{i}$ divided by a set factor $\lambda_{s}$:
$\mathbf{S}_{j}=\frac{1}{\lambda_{s}}\mathbf{S}_{i}$ (4)
$\mathbf{R}_{j}=\mathbf{R}_{i},\ SH_{j}=SH_{i},\ o_{j}=o_{i}$
$\boldsymbol{\mu}_{j}\sim\mathcal{N}(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$
where $\boldsymbol{\Sigma}_{i}$ is the covariance matrix of $G_{i}$. Then, to tackle the density increase in the overlap of the two Gaussians, we reduce the opacity of the original Gaussian. This is accomplished by multiplying its opacity by a specified factor $\beta$:
$o^{\prime}_{i}=\beta o_{i}$ (5)
Figure 4: Intermediate results after applying residual split. We saved the Gaussians at 3,000 and 6,000 iterations after applying residual split. As observed, residual split can adaptively handle over-reconstruction and under-reconstruction by refining blurry regions with finer Gaussians while also effectively reconstructing missing structures. Zoom in for better view.
Unlike clone and split which lack adaptiveness, residual split can fill in missing geometry and recover details adaptively without requiring a trade-off, as shown in Fig. 2 (c). We further illustrate this using intermediate results, as shown in Fig. 4. In areas where details are challenging to recover, down-scaled Gaussians of finer granularity are progressively introduced to refine and reveal intricate details. On the other hand, since densification does not change the scale of the original Gaussian, the total coverage is increased after densification, helping to reconstruct missing structures. Moreover, our method’s nonidentical approach with hierarchical Gaussians of varying scales enables efficient shape fitting. As illustrated by the blue rectangles in Fig. 1, this hierarchy allows us to fill textureless regions with a few large Gaussians, significantly reducing redundancy. Consequently, our approach delivers higher rendering quality while using fewer Gaussians.
3.2.2 Image Pyramid as Supervision
Building on our motivation in Sec. 1, we designed a coarse-to-fine supervision schedule that gradually integrates information into the scene. Specifically, based on the original multi-view training pictures, we build an image pyramid $\{IP_{i}\}_{i=1}^{L}$ consisting of $L$ layers. $IP_{i}$ corresponds to the set of images at different viewpoints in layer $i$. $IP_{L}$ is the set of original pictures:
$IP_{i}=\{P^{v}_{i}\}_{v\in V}$ (6)
where $V$ is the set of viewpoints and $P^{v}_{i}$ refers to the picture at viewpoint $v$ in the $i^{\text{th}}$ layer. Let $(H_{L}^{v},W_{L}^{v})$ be the resolution of $P_{L}^{v}$, and $(H_{i}^{v},W_{i}^{v})$ represent the resolution of $P_{i}^{v}$. Then there is:
$(H_{i}^{v},W_{i}^{v})=(H_{L}^{v}/2^{L-i},W_{L}^{v}/2^{L-i})$ (7)
Correspondingly, our training process is divided into $L$ stages. At stage $i$, we train our model under the supervision of $IP_{i}$, increasing the resolution of the training pictures throughout the stages. Due to the variation of frequency bandwidth in different levels in the image pyramid, high-frequency information is gradually added to the scene. Thus, we can force our model to learn coarse and low-frequency features at early stages and construct an overall structure of the whole scene. At later stages, the focus shifts to capturing high-frequency components, carving details above the coarse expression.
3.2.3 Varying Gradient Threshold
To further emphasize progressive training, an incremental approach for selecting Gaussians to densify is designed. Specifically, as iterations progress, the model’s focus gradually shifts toward finer details. To support this, we aim to encourage the densification of coarse Gaussians to capture more detail over time. Due to the nonidentical of our residual split, we can expect the generated Gaussians to be finer than the original ones. This allows us to track the relative fineness level of Gaussians, denoted as $l_{i}$. We assign $l_{i}$ of initial Gaussians to 0. If the level of the Gaussian being densified is $l_{i}$, we assign the newly added Gaussian a level of $l_{i}+1$. Thus, as $l_{i}$ gets higher, Gaussians tend to be smaller and finer. Note that $l_{i}$ is only utilized during training and discarded afterward.
A single stage is then divided evenly into $K$ substages, creating $L\times K$ substages in total. The gradually increasing substage can be seen as an indicator of the scene details level. At substage $k$, we promote additional densification for Gaussians with $l_{i}<k$ and prioritize those at lower levels, achieved by applying varying gradient thresholds. Specifically, the threshold $\tau_{k,i}$ for a Gaussian $G_{i}$ is set as:
$\tau_{k,i}=\left\{\begin{aligned} &\tau,l_{i}\geq k\\ &\frac{\tau}{\alpha^{k-l_{i}}},l_{i}<k\end{aligned}\right.$ (8)
where $\tau$ is a set threshold and $\alpha>1$ is a chosen value.
4 Experiments
Table 1: Quantitative comparison to previous methods on real-world datasets. Our method achieves SOTA in the Mip-NeRF360 dataset and comparable results with SOTA on the Tanks&Temples and Deep Blending datasets. indicates results generated by retrained models for different rendering resolution.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem.
Plenoxels 23.08 0.626 0.463 2.1GB 21.08 0.719 0.379 2.3GB 23.06 0.795 0.510 2.7GB
Instant-NGP 25.59 0.699 0.331 48MB 21.92 0.745 0.305 48MB 24.96 0.817 0.390 48MB
Mip-NeRF360 27.69 0.792 0.237 8.6MB 22.22 0.759 0.257 8.6MB 29.40 0.901 0.245 8.6MB
Scaffold-GS 27.69 0.812 0.225 176MB 23.96 0.853 0.177 87MB 30.21 0.906 0.254 66MB
Octree-GS 27.52 0.812 0.218 124MB 24.52 0.866 0.153 84MB 30.41 0.913 0.238 93MB
3D-GS 27.21 0.815 0.214 734MB 23.14 0.841 0.183 411MB 29.41 0.903 0.243 676MB
AbsGS 27.49 0.820 0.191 728MB 23.73 0.853 0.162 304MB 29.67 0.902 0.236  [^4] MB
Pixel-GS 27.52 0.822 0.191 1.32GB 23.78 0.853 0.151 1.06GB 28.91 0.892 0.250 1.09GB
FreGS 27.85 0.826 0.209 - 23.96 0.849 0.178 - 29.93 0.904 0.240 -
Mini-Splatting-D 27.51 0.831 0.176 1.11GB 23.23 0.853 0.140 1.01GB 29.88 0.906 0.211 1.09GB
Ours-Small (AbsGS) 27.94 0.830 0.191 342MB 24.33 0.862 0.150 187MB 30.01 0.906 0.234 289MB
Ours (3D-GS) 28.00 0.831 0.187 600MB 24.26 0.865 0.141 354MB 29.93 0.903 0.232 596MB
Ours (AbsGS) 28.00 0.833 0.174 698MB 24.38 0.867 0.132 351MB 29.91 0.902 0.227 586MB
4.1 Experiment Settings
Following previous works [24, 43, 48, 29, 35, 13], evaluation of our model was done on three datasets, including all nine scenes from Mip-NeRF360 [1], two scenes from Tanks&Temples [25], and two scenes from DeepBlending [18]. The test and train splits follow those of the original 3D-GS [24], and the image resolution is also maintained as in 3D-GS.
Three commonly used metrics were used to quantify the fidelity of rendered images: PSNR, SSIM [38], and LPIPS [49]. We also report the memory consumption as an indicator of the number of Gaussians.
We trained three versions of our model: two with AbsGS [43], leveraging its enhanced indication of Gaussians for densification—one standard version and one small version with higher gradient thresholds—and a third version using 3D-GS as the baseline. We applied the periodic opacity reduction from AbsGS across all three versions, as it proved to be a more effective alternative to the opacity reset used in 3D-GS. Additionally, we extended this reduction beyond the densification phase, finding it further minimized redundancy without compromising fidelity. Following all 3D-GS-like methods, we trained our model for 30k iterations. The densification stops at 12000 iterations. For the version with AbsGS, $\tau$ was set to $0.00067$, the small version to $0.0016$; for the version with 3D-GS, $\tau$ was set to $0.00028$. $L$ was set to $3$ and $K$ was set to $3$. Our first stage lasts for 2500 iterations, the second stage for 3500 iterations, and the rest iterations are in the last stage. A 500-iteration warm-up phase was introduced between stages with densification disabled. $\alpha$ was set to $2^{\frac{1}{3}}$, $\lambda_{s}$ was set to $1.6$ and $\beta$ was set to $0.3$. The loss function remains the same as in 3D-GS [24], consisting of a combination of $\mathcal{L}_{1}$ and D-SSIM losses. All experiments were done on an NVIDIA RTX3090 GPU.
4.2 Comparisons
To assess the effectiveness of our method, we compared it with several 3D-GS variants: 3D-GS [24], Scaffold-GS [29], Octree-GS [35], FreGS [48], AbsGS [43], Pixel-GS [50] and Mini-Splatting-D [13]. Note that Scaffold-GS and Octree-GS use a latent representation rather than the original 3D-GS representation. Some radiance-field-based methods are also selected for comparison, including Plenoxels [45], Instant-NGP [32] and Mip-NeRF360 [1]. We found that the resolution of Mip-NeRF360 dataset images varies across different works. Following the resolution setup in the 3D-GS work, we retrain some models on the Mip-NeRF360 dataset. Additionally, for works that did not report their storage, we obtained and reported it based on their released code.
Quality Comparisons.
In Table 1, we present quantitative results comparing the rendering quality of our method with other approaches. From the table, we can see that our method achieves SOTA performance across all metrics on the Mip-NeRF360 [1] dataset. On the Tanks&Temples [25] dataset, we also achieve the best SSIM and LPIPS metrics, which closely reflect human perception. In the Deep Blending [18] dataset, while our method does not achieve the highest score, it shows a notable improvement over the baseline. Additionally, our model variant without AbsGS [43] maintains similar performance to the full model, with only a slight reduction in LPIPS. Notably, our smaller model version achieves strong performance with reduced memory usage, showcasing the efficiency and scalability of our approach. In Fig. 5, we present qualitative comparisons between our method and several 3D-GS variants [24, 50, 43] across various scenes, including both indoor and outdoor settings. It is evident that our model captures finer details, as shown in the second, third, and fourth rows, constructs missing geometry more effectively (last row), and achieves more accurate geometry reconstruction (first row).
Figure 5: Qualitative comparisons of ResGS with three 3D-GS variations on a variety of indoor and outdoor scenes. Our approach captures more intrinsic details and acquires more complete geometry in complex scenes.
Table 2: Efficiency comparison with former methods.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
FPS Time FPS Time FPS Time
Octree-GS 119 29m55s 126 23m8s 151 34m38s
Scaffold-GS 103 23m53s 137 15m37s 141 19m40s
3D-GS 103 27m16s 127 15m34s 101 24m48s
AbsGS 132 28m48s 193 13m35s 185 22m3s
Mini-Splatting-D 74 34m42s 75 27m39s 100 30m37s
Pixel-GS 65 40m6s 70 29m29s 70 34m1s
Ours-Small (AbsGS) 141 20m56s 206 11m35s 202 19m26s
Ours (AbsGS) 100 29m24s 150 16m29s 153 23m33s
Table 3: Analyzing residual split upon multiple 3D-GS variants. The performance consistently increases on these pipelines, indicating that our Residual split is a better densification operation. 3D-GS indicates a retrained version of 3D-GS [24] for better performance.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem
3D-GS 27.21 0.815 0.214 734MB 23.14 0.841 0.183 411MB 29.41 0.903 0.243 676MB
3D-GS 27.38 0.813 0.215 821MB 23.66 0.844 0.179 434MB 29.48 0.900 0.247 663MB
3D-GS + residual split 27.44 0.813 0.216 586MB 23.76 0.845 0.178 318MB 29.75 0.902 0.240 546MB
AbsGS 27.49 0.820 0.191 728MB 23.73 0.853 0.162 304MB 29.67 0.902 0.236  [^4] MB
AbsGS + residual split 27.71 0.827 0.185 712MB 24.05 0.857 0.161 285MB 29.68 0.905 0.234 421MB
Pixel-GS 27.52 0.822 0.191 1.32GB 23.78 0.853 0.151 1.06GB 28.91 0.892 0.250 1.09GB
Pixel-GS + residual split 27.62 0.821 0.191 1.00GB 24.05 0.856 0.148 754MB 29.75 0.899 0.235 914MB
Mini-Splatting-D 27.51 0.831 0.176 1.11GB 23.23 0.853 0.140 1.01GB 29.88 0.906 0.211 1.09GB
Mini-Splatting-D + residual split 27.64 0.831 0.176 1.04GB 23.49 0.853 0.140 967MB 29.94 0.903 0.210 1.02GB
Efficiency Comparisons.
Here, we provide the computational costs of our method in comparison with other approaches. The results are shown in Table. 2. From the table, we can see that our method involves a slight increase in training time while delivering much higher rendering quality than methods with lower rendering quality like 3D-GS [24] and AbsGS [43]. Our method is also significantly more efficient than other methods like Mini-Splatting-D [13], with comparable rendering quality. Notably, our small model provides the fastest FPS and shortest training time in all compared methods while maintaining high fidelity.
4.3 Analysis
4.3.1 Residual Split
Compatibility.
We analyze the compatibility of our novel densification operation by replacing split and clone in multiple 3D-GS variants [24, 43, 50, 13] with residual split. The hyperparameters were kept the same as in the original pipelines. The results are shown in Table 3. The table shows that our method leads to a notable increase in PSNR metrics across all pipelines, while for other metrics, some exhibit improvements, and others remain approximately the same. Additionally, memory consumption is reduced after incorporating residual split compared to the original pipeline. This underscores our method’s ability to improve rendering quality and reduce redundancy, while also demonstrating its compatibility and potential for broader application across 3D-GS-oriented works as an effective alternative to split and clone.
Figure 6: Impact of the opacity reduction weight $\beta$ for residual split.
Figure 7: Impact of the scale reduction factor $\lambda_{s}$ for residual split.
Hyperparameter Sensitivities.
Since residual split involves some hyperparameters: opacity reduction weight $\beta$ and scale reduction factor $\lambda_{s}$, we conducted a corresponding sensitivity analysis. We tested our method upon the Tanks&Temples dataset while applying different $\beta\in[0.05,1.0)$ and $\lambda_{s}\in[1.05,2.0]$. All the models are trained with similar storage. The LPIPS scores are shown in Fig. 6 and 7. As shown in the figures, LPIPS scores remain relatively stable when $\beta<0.4$ and $\lambda_{s}>1.55$. These results indicate that our method is robust to hyperparameter variations, with only a 0.006 difference in LPIPS between the best and worst cases, and a relatively broad range where rendering quality remains optimal.
4.3.2 Ablation Studies
Table 4: Ablation Study on our pipeline. Here, IP denotes image pyramid supervision, RS represents using residual split for densification, and VT indicates applying varying gradient threshold. Each component contributes to our method’s effectiveness.
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Base 27.41 0.817 0.189
Base + RS 27.66 0.825 0.183
Base + IP 27.54 0.823 0.182
Base + RS + IP 27.88 0.831 0.178
Base + RS + IP + VT (full) 28.00 0.833 0.174
To assess the effectiveness of individual modules in our proposed pipeline—residual split, image pyramid as supervision, and varying gradient threshold—ablation studies on the Mip-NeRF360 [1] dataset were conducted. For the varying gradient threshold, we only conducted studies on its removal, as it is highly interdependent with the other two modules. The results are shown in Table 4. It is evident that both residual split and the image pyramid supervision individually enhance the quality of our baseline. However, we observe that adding image pyramid supervision alone does not lead to significant improvements. This is because the split-and-clone operations do not perform well within a coarse-to-fine schedule. In the early stages, when the overall structure is still being constructed, split is primarily chosen, which exacerbates the issue rather than improving it. On the other hand, in later iterations, when fine details need to be recovered, clone is selected and does not address the issue properly. Conversely, residual split works perfectly due to its adaptiveness. This is evident from the notable improvement of Base + RS + IP over both Base + IP and Base + RS, which also shows that a coarse-to-fine schedule supplements residual split as expected. Additionally, applying the varying gradient threshold further boosts performance. In conclusion, each module contributes to improved results.
4.4 Discussions and Limitations
Our results show that our method significantly improves rendering quality on the Mip-NeRF360 [1] and Tanks&Temples [25] datasets. However, for Deep Blending [18], our method did not yield a significant boost in performance. This variation is primarily due to the challenges specific to each dataset: Mip-NeRF360 and Tanks&Temples require capturing fine textures and recovering missing geometry, whereas Deep Blending focuses on handling occlusions and reconstructing accurate geometry in weakly textured areas. Our method mainly contributes to retrieving fine details and sufficient geometry and does not address the issue of occlusion and weak-textured areas, suggesting it to be a topic for future optimization.
5 Conclusion
In this paper, we analyzed the limitations of the current 3D-GS densification operation. To address these issues, we propose residual split, a residual empowered densification operation that adaptively handles both over-reconstruction and under-reconstruction. Additionally, we introduce ResGS, a coarse-to-fine 3D-GS pipeline to supplement residual split. Specifically, we use an image pyramid for supervision to gradually recover the overall structure to fine details during Gaussian optimization. We also design a Gaussian selection scheme to further densify coarse Gaussians progressively. Experiments show our pipeline achieves SOTA rendering quality while residual split is highly compatible with other 3D-GS variants.
Supplementary Material
6 Additional Experiment Results
In this section, we provide additional experimental results of our work.
6.1 Variance of PSNR on Deep Blending Dataset
A notable variance in the PSNR metrics of our method is observed on the Deep Blending [18] dataset. We tested our method on the Deep Blending dataset 10 times and reported the average performance and population standard deviation. The results are shown in Table 5. The table demonstrates that the PSNR scores on the Deep Blending dataset exhibit a notable variance. The underlying reason, as discussed in our limitations, is that our method does not address issues related to occlusion and overfitting, making the Deep Blending dataset particularly challenging for our approach. This highlights opportunities for future optimization and improvement. Furthermore, we report the average performance and population standard deviation over 10 runs for the other two datasets, Mip-NeRF360 [1] and Tanks&Temples [25], as presented in Table 6 and Table 7, respectively. The PSNR metrics for these two datasets do not show a notable variance, showcasing the robustness of our method in most scenarios.
Table 5: Average and standard deviation results of three metrics on the Deep Blending dataset.
Dataset Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Avg Stdev Avg Stdev Avg Stdev
Ours-Small (AbsGS) 29.71 0.211 0.903 0.002 0.235 0.002
Ours (3D-GS) 29.64 0.205 0.900 0.003 0.233 0.002
Ours (AbsGS) 29.68 0.173 0.900 0.002 0.228 0.002
Table 6: Average and standard deviation results of three metrics on the Mip-NeRF360 dataset.
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Avg Stdev Avg Stdev Avg Stdev
Ours-Small (AbsGS) 27.93 0.010 0.830 0.001 0.191 0.001
Ours (3D-GS) 27.99 0.011 0.831 0.001 0.187 0.001
Ours (AbsGS) 27.99 0.011 0.833 0.001 0.174 0.001
Table 7: Average and standard deviation results of three metrics on the Tank&Temples dataset.
Dataset Tank&Temples
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Avg Stdev Avg Stdev Avg Stdev
Ours-Small (AbsGS) 24.21 0.081 0.862 0.001 0.151 0.001
Ours (3D-GS) 24.26 0.058 0.864 0.001 0.141 0.001
Ours (AbsGS) 24.27 0.063 0.866 0.001 0.133 0.001
6.2 Further Analysis of Residual Split on 3D-GS
Table 8: Results of residual split upon 3D-GS with varied gradient threshold.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem
3D-GS 27.21 0.815 0.214 734MB 23.14 0.841 0.183 411MB 29.41 0.903 0.243 676MB
3D-GS + residual split 27.48 0.815 0.213 722MB 23.85 0.846 0.174 374MB 29.83 0.901 0.238 640MB
In our paper, we tested the compatibility of our residual split across multiple pipelines without modifying any hyper-parameters. Here, we present the results of applying residual split to 3D-GS [24] while adjusting the densification threshold $\tau$ to maintain the same storage size as 3D-GS, further demonstrating the capability of our method to enhance rendering quality. The results are shown in Table 8, showing that our method yields a further performance increase.
Table 9: Results of our method without extended periodic opacity reduction. Here, No EPOR refers to our method trained without applying the extended periodeic opacity reduction. No EPOR + CT refers to the model without the extended periodic opacity reduction but with an adjusted gradient threshold to ensure that the final storage size is approximately the same as our full model with extended periodic opacity reduction.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem
Ours (No EPOR) 28.02 0.833 0.173 973MB 24.32 0.867 0.130 591MB 29.86 0.902 0.227 911MB
Ours (No EPOR + CT) 28.00 0.833 0.181 608MB 24.35 0.866 0.139 380MB 29.96 0.904 0.230 564MB
Ours (full) 28.00 0.833 0.174 698MB 24.38 0.867 0.132 351MB 29.91 0.902 0.227 586MB
6.3 Periodic Opacity Reduction
3D-GS [24] often encounters redundancy issues, which we categorize into two main causes: the first is the use of an excessive number of small-scale Gaussians to represent coarse areas, and the second is the overlap between Gaussians, resulting in redundant Gaussians that contribute minimally to the rendered image. Our residual split can solve the first issue but does not address the second. To tackle the second issue, as mentioned in our experiment settings, we implemented the periodic opacity reduction technique in [43, 5]. Specifically, the opacity of Gaussians is periodically reduced, causing the opacity of those that contribute little to the rendered image to diminish to a negligible value, allowing them to be effectively pruned. Additionally, we extended this operation to occur after the densification stage, discovering that it further reduces redundancy without compromising fidelity. Here, we present the experimental results of our method without extending the periodic opacity reduction operation, as shown in Table 9. From the table, we can observe that extending the periodic opacity reduction operation effectively reduces redundancy without compromising the fidelity of the rendered images. Furthermore, when the operation is not extended and the gradient threshold is adjusted to maintain storage approximately equal to our full model, the rendering quality shows only a slight decrease in LPIPS metrics. This indicates that while extending the operation can enhance the performance of our method, its overall impact on the performance improvement is relatively minor.
6.4 Image Resolution of Mip-NeRF360 Dataset
In our paper, we mentioned that the image resolution on the Mip-NeRF360 [1] dataset varies across different works. 3D-GS [24] uses the provided “images_4”, the official downscaled 4 times images from the dataset, for outdoor scenes and “images_2” for indoor scenes. Most works [43, 13, 48] follow the 3D-GS setting, but Scaffold-GS [29], Octree-GS [35], Pixel-GS [50] uses different settings. Specifically, Scaffold-GS and Octree-GS downscale original images to 1.6K resolution, while Pixel-GS downscales original images by a factor of 4 for outdoor scenes and 2 for indoor scenes during the training process, instead of using the provided downscaled images. This discrepancy results in differences in experimental settings. To further demonstrate the performance of our method, we provide its results under the two different settings, as shown in Table 10 and 11. The tables show that our method prevails even under these alternate settings.
Table 10: Results on the Mip-NeRF360 dataset with same setting as Scaffold-GS and Octree-GS.
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Scaffold-GS 27.71 0.813 0.221
Octree-GS 27.73 0.815 0.217
Ours-Small (AbsGS) 28.04 0.831 0.191
Ours (3D-GS) 28.10 0.832 0.187
Ours (AbsGS) 28.14 0.834 0.174
Table 11: Results on the Mip-NeRF360 dataset with same setting as Pixel-GS.
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Pixel-GS 27.88 0.834 0.176
Ours-Small (AbsGS) 28.21 0.841 0.176
Ours (3D-GS) 28.28 0.843 0.172
Ours (AbsGS) 28.30 0.845 0.160
6.5 Per Scene Results
We present the per-scene results of the used metrics in Table 12-17. For works that did not report per-scene results, we obtained and reported them using their released code. The works selected for comparison are mainly the same as in the paper: 3D-GS [24], Scaffold-GS [29], Octree-GS [35], AbsGS [43], Pixel-GS [50], Mini-Splatting-D [13], Plenoxels [45], Instant-NGP [32] and Mip-NeRF360 [1]. Note that we did not show the results of FreGS [48] since they did not report their per-scene results or release their code. Works retrained on the Mip-NeRF360 [1] are marked with a .
Table 12: Per-scene PSNR metrics of Mip-NeRF360 dataset.
Method Scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Plenoxels 21.91 20.10 23.49 20.66 22.25 27.59 23.62 23.42 24.67
Instant-NGP 22.17 20.65 25.07 23.47 22.37 29.69 26.69 29.48 30.69
Mip-NeRF360 24.31 21.65 26.88 26.18 22.93 31.47 29.45 31.99 33.40
Scaffold-GS 25.07 21.36 27.35 26.64 22.93 32.04 29.50 31.63 32.71
Octree-GS 25.05 21.33 27.58 26.41 22.81 32.09 29.48 30.89 31.71
3D-GS 25.25 21.52 27.41 26.55 22.49 30.63 28.70 30.32 31.98
AbsGS 25.29 21.35 27.48 26.71 21.99 31.61 29.03 31.62 32.32
Pixel-GS 25.21 21.49 27.42 26.84 22.09 31.45 29.05 31.65 32.53
Mini-Splatting-D 25.55 21.50 27.67 27.11 22.13 31.41 28.72 31.75 31.72
Ours-Small (AbsGS) 25.62 21.80 27.72 27.27 22.90 32.41 29.32 31.95 32.48
Ours (3D-GS) 25.55 22.07 27.71 27.12 22.82 32.48 29.41 32.09 32.74
Ours (AbsGS) 25.62 21.78 27.82 27.19 22.57 32.51 29.50 32.26 32.78
Table 13: Per-scene SSIM metrics of Mip-NeRF360 dataset.
Method Scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Plenoxels 0.496 0.431 0.6063 0.523 0.509 0.8417 0.759 0.648 0.814
Instant-NGP 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906
Mip-NeRF360 0.685 0.584 0.809 0.745 0.631 0.910 0.892 0.917 0.938
Scaffold-GS 0.755 0.590 0.858 0.765 0.639 0.923 0.910 0.926 0.943
Octree-GS 0.759 0.597 0.864 0.763 0.643 0.924 0.907 0.916 0.930
3D-GS 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938
AbsGS 0.783 0.623 0.871 0.780 0.617 0.925 0.911 0.929 0.945
Pixel-GS 0.775 0.633 0.867 0.784 0.630 0.920 0.911 0.929 0.944
Mini-Splatting-D 0.798 0.642 0.878 0.804 0.640 0.928 0.913 0.934 0.946
Ours-Small (AbsGS) 0.792 0.634 0.872 0.802 0.652 0.929 0.914 0.932 0.945
Ours (3D-GS) 0.790 0.639 0.874 0.800 0.650 0.929 0.916 0.934 0.947
Ours (AbsGS) 0.797 0.647 0.876 0.804 0.645 0.931 0.918 0.934 0.948
Table 14: Per-scene LPIPS metrics of Mip-NeRF360 dataset.
Method Scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Plenoxels 0.506 0.521 0.3864 0.503 0.540 0.4186 0.441 0.447 0.398
Instant-NGP 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205
Mip-NeRF360 0.301 0.344 0.170 0.261 0.339 0.211 0.204 0.127 0.176
Scaffold-GS 0.233 0.352 0.120 0.236 0.331 0.216 0.203 0.130 0.207
Octree-GS 0.225 0.340 0.108 0.233 0.310 0.203 0.202 0.141 0.219
3D-GS 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205
AbsGS 0.171 0.270 0.100 0.195 0.278 0.200 0.189 0.121 0.190
Pixel-GS 0.182 0.263 0.100 0.186 0.280 0.213 0.185 0.120 0.193
Mini-Splatting-D 0.158 0.255 0.090 0.169 0.262 0.190 0.172 0.114 0.175
Ours-Small (AbsGS) 0.175 0.290 0.099 0.188 0.279 0.196 0.184 0.117 0.189
Ours (3D-GS) 0.168 0.295 0.092 0.179 0.280 0.192 0.177 0.114 0.185
Ours (AbsGS) 0.156 0.251 0.089 0.170 0.250 0.187 0.172 0.112 0.179
Table 15: Per-scene PSNR metrics of Tank&Temples and Deep Blending dataset.
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 23.22 18.93 23.14 22.98
Instant-NGP 23.38 20.46 28.26 21.67
Mip-NeRF360 24.91 19.52 29.14 29.66
Scaffold-GS 25.77 22.15 29.80 30.62
Octree-GS 26.27 22.77 29.87 30.95
3D-GS 25.19 21.10 28.77 30.04
AbsGS 25.74 21.72 29.20 30.14
Pixel-GS 25.49 22.13 28.02 29.79
Mini-Splatting-D 25.43 21.04 29.32 30.43
Ours-Small (AbsGS) 26.07 22.40 29.64 30.38
Ours (3D-GS) 26.08 22.65 29.54 30.32
Ours (AbsGS) 26.14 22.61 29.51 30.32
Table 16: Per-scene SSIM metrics of Tank&Temples and Deep Blending dataset.
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 0.774 0.663 0.787 0.802
Instant-NGP 0.800 0.689 0.854 0.779
Mip-NeRF360 0.857 0.660 0.901 0.900
Scaffold-GS 0.883 0.822 0.907 0.904
Octree-GS 0.896 0.835 0.912 0.914
3D-GS 0.879 0.802 0.899 0.906
AbsGS 0.888 0.818 0.898 0.907
Pixel-GS 0.883 0.823 0.885 0.899
Mini-Splatting-D 0.890 0.817 0.905 0.908
Ours-Small (AbsGS) 0.889 0.834 0.905 0.907
Ours (3D-GS) 0.890 0.840 0.901 0.905
Ours (AbsGS) 0.893 0.841 0.900 0.905
Table 17: Per-scene LPIPS metrics of Tank&Temples and Deep Blending dataset.
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 0.335 0.422 0.521 0.499
Instant-NGP 0.249 0.360 0.352 0.428
Mip-NeRF360 0.159 0.354 0.237 0.252
Scaffold-GS 0.147 0.206 0.250 0.258
Octree-GS 0.118 0.198 0.231 0.244
3D-GS 0.148 0.218 0.244 0.241
AbsGS 0.132 0.193 0.241 0.233
Pixel-GS 0.122 0.180 0.258 0.243
Mini-Splatting-D 0.100 0.181 0.218 0.204
Ours-Small (AbsGS) 0.124 0.177 0.233 0.235
Ours (3D-GS) 0.118 0.164 0.232 0.231
Ours (AbsGS) 0.106 0.159 0.231 0.223