### v3

# ResGS: Residual Densification of 3D Gaussian for Efficient Detail Recovery

Abstract
Recently, 3D Gaussian Splatting (3D-GS) has prevailed in novel view synthesis, achieving high fidelity and efficiency. However, it often struggles to capture rich details and complete geometry. Our analysis reveals that the 3D-GS densification operation lacks adaptiveness and faces a dilemma between geometry coverage and detail recovery. To address this, we introduce a novel densification operation, residual split, which adds a downscaled Gaussian as a residual. Our approach is capable of adaptively retrieving details and complementing missing geometry. To further support this method, we propose a pipeline named ResGS. Specifically, we integrate a Gaussian image pyramid for progressive supervision and implement a selection scheme that prioritizes the densification of coarse Gaussians over time. Extensive experiments demonstrate that our method achieves SOTA rendering quality. Consistent performance improvements can be achieved by applying our residual split on various 3D-GS variants, underscoring its versatility and potential for broader application in 3D-GS-based applications. Project page: https://yanzhelyu.github.io/resgs.github.io/.
Figure 1: We propose ResGS, a pipeline to boost the rendering quality of 3D-GS [23] while improving efficiency. (a) The 3D-GS densification operations, split and clone, rely on binary selection and cannot adaptively address the challenges they encounter. (b) Our ResGS employs a single densification operation, residual split, that adaptively addresses under-reconstruction and over-reconstruction. (c) 3D-GS tends to generate Gaussians with sizes in small variation and low rendering quality. (d) From the same initialization, our method can capture fine details and retrieve sufficient geometry effectively while reducing redundancy.
[…]
1 Introduction
Figure 2: An illustration of the limitation of split and clone. (a) selects between split and clone based on a high threshold, (b) a low threshold, and in (c) the densification operation is replaced as residual split. We provide visualizations of the final rendered image. A high threshold ensures sufficient structural coverage but causes over-reconstruction, leading to blurry rendering. Conversely, a low threshold mitigates blurry rendering but results in under-reconstruction, compromising coverage and creating a difficult trade-off. Our method, however, adaptively tackles both issues without facing this trade-off.
Novel View Synthesis (NVS) has continued to be a hot topic throughout the years. Traditional methods like point clouds and meshes [4, 13, 9] can achieve high rendering speed but often lack fidelity. On the other hand, neural volumetric approaches [29, 1, 30] can achieve high fidelity for novel views but significantly sacrifice rendering efficiency. Recently, 3D-GS [23] has shown both high fidelity and fast rendering speed for novel view synthesis. However, 3D-GS still suffers from setbacks, often failing to capture fine details or recover complete geometry, as shown by the red and orange rectangles in Fig. 1 and Fig. 2.
While various approaches [48, 50, 43, 12] have been proposed to tackle these challenges and have significantly advanced the rendering quality of 3D-GS-based methods to state-of-the-art (SOTA) levels, they seldom adapt the densification operation of 3D-GS. In this paper, we go through the densification procedure and propose a novel adaptive densification approach to improve reconstruction quality.
The densification operation used in 3D-GS-based methods consists of two operations: split and clone. The split operation is designed to address over-reconstruction, where a single Gaussian covers too much area, and the clone operation aims to compensate for under-reconstruction. However, selecting between the two operations properly in different areas is challenging and existing methods rely on a threshold-based decision mechanism. As illustrated in Fig. 1 (a), Gaussians whose scale $s=max(s_{x},s_{y},s_{z})$ is larger than a predefined threshold $\tau_{s}$ undergo split, while those below it use clone. The simplicity of this threshold-based densification usually leads to suboptimal results and encounters a dilemma in threshold selection. As highlighted by the red rectangles in Fig. 2 (a), a high threshold results in insufficient split preventing the model from capturing fine details and leading to blurry rendering. On the other hand, a low threshold suppresses clone operations, leading to missing geometry, as illustrated in the orange boxes in Fig. 2 (b). Furthermore, 3D-GS tends to first split and then clone, resulting in objects being fitted with Gaussians of similar scales. As highlighted by the blue rectangles in Fig. 1, 3D-GS fits textureless areas with an excessive number of small Gaussians, which is redundant.
To tackle the aforementioned problems, we propose residual split, a novel densification operation in 3D Gaussian splatting. As illustrated in Fig. 1 (b), the core idea is to spawn a downscaled replica as a residual component while reducing the original Gaussian’s opacity for any under-optimized Gaussian. This simultaneously expands spatial coverage and injects finer details, eliminating the dilemma of selecting split or clone.
However, directly applying residual split does not produce significant improvements. This is due to the inefficiency the 3D-GS initialization. At early iterations, the scene is represented by a few coarse Gaussians that very likely face both over-reconstruction and under-reconstruction. Gaussians added through densification are challenging to optimize, as they must address both issues simultaneously. It’s desired to develop an approach that decouples coverage and detail optimization.
Correspondingly, we propose a coarse-to-fine training pipeline, ResGS, that progressively shifts the focus of training, facilitating optimization. Firstly, we structure the training process into stages, employing multi-scale images for gradual supervision, ensuring that the early stages focus on the overall structure and the later stages focus on details. Secondly, we develop a progressive Gaussian selection scheme for densification, which gradually encourages coarser Gaussians to undergo further refinement. Therefore, finer structures are introduced at later iterations, avoiding premature optimization. These two techniques simplify optimization and thereby improve rendering quality.
Our contributions are summarized as follows:
  • We propose a novel densification operation for 3D Gaussians based on a residual strategy to overcome the limitations of fixed threshold for determining split or clone in 3D-GS;
  • We develop a coarse-to-fine training pipeline, ResGS. Equipped with the proposed residual split, it incorporates a multi-scale supervision with image pyramids and a Gaussian selection scheme to enhance rendering quality;
  • Our method achieves SOTA rendering quality in NVS on three datasets, including Mip-NeRF360 [1], Tanks&Temples [24], and DeepBlending [17]. Furthermore, applying our proposed residual split to multiple 3D-GS variants consistently brings performance improvements, demonstrating that our new densification operation is more effective than densification with a fixed scale threshold.
2 Related Work
2.1 Novel View Synthesis
Novel view synthesis (NVS) aims to acquire novel view images for a scene through a set of multi-view images. As a pioneer, NeRF [29] introduced an MLP-based implicit representation. More specifically, it uses an MLP to predict the color and opacity of a point in a scene and uses differentiable volume rendering to render images. Due to its photo-realistic quality, it has been extensively studied [7, 21, 6, 47, 26, 8, 1, 2]. However, NeRF-based methods often suffer from long training and rendering time. Therefore, many works [31, 32, 18, 19] have been proposed to extend NeRF into real-time rendering.
Recently, 3D-GS [23] has stirred an evolution in the field by achieving both high-fidelity and real-time rendering. Due to its high performance, many works have enabled 3D-GS into many downstream tasks such as SLAM [28, 41], 3D generation [10, 37, 44, 52], 4D scene modeling [42, 39, 15, 25], and surface reconstruction [20, 14].
However, the original 3D-GS still has limitations, and myriad works have been proposed to increase the render quality and efficiency of 3D-GS. Mip-Splatting [46] introduced a 2D dilation filter and a 3D smoothing filter to mitigate aliasing and artifacts caused by changes in focal length. Scaffold-GS [27] proposed the concept of neural Gaussians which is to tie Gaussians around an anchor and use an MLP to predict its attributes. Thus, the model is more robust to significant view changes and can improve rendering quality. AbsGS [43] discovered that when accumulating the view space gradients used in densification through pixels, there could be a collision, for gradients at various pixels might have different positivities. They applied an absolute function before summing through pixels to mitigate the issue. Pixel-GS [50] analyzed 3D-GS artifacts from the pixel perspective and introduced a method to select Gaussians for densification based on this analysis. GaussianPro [11] introduced a progressive propagation strategy to propagate Gaussians to under-initialized areas. To reduce redundancy and increase rendering quality, Mini-Splatting [12] proposed a densification and simplification algorithm to reorganize Gaussians’ spatial positions.
With the emergence of numerous works focused on optimizing NeRF and 3D-GS, progressive training schedules—proven effective in multiple scenarios [3, 16, 34, 22] have earned attention.
2.2 Progressive Training in NVS
Several progressive training schedules have been explored upon NeRF-based methods. BungeeNeRF [40] captured different levels of details in large city-scale datasets by gradually adding closer views to the training views and appending residual blocks to the MLP throughout the training process. Pyramid NeRF [51] employs a coarse-to-fine strategy, using an image pyramid to progressively introduce high-frequency details while simultaneously subdividing the scene representation.
Recently, with the prevalence of 3D-GS, many progressive training pipelines have been proposed. Octree-GS [33] proposed an Octree-based 3D-GS approach, and the Octree depth is progressively deepened. FreGS [48] designed a progressive frequency regulation for 3D-GS which is to calculate the loss of rendered images in the frequency domain and gradually add higher frequency components.
3 Our Method
Figure 3: Overview of our ResGS pipeline. (a) The core of our pipeline, residual split, involves adding a downscaled replicate and then reducing the opacity of the original Gaussian. (b) We assign initial Gaussians a temporary attribute $l_{i}=0$ for densification selection, which is discarded after training. Next, the pipeline is split into $L$ stages, with each stage trained on images downscaled using an image pyramid. Each single stage is further divided evenly into $K$ substages, for selecting Gaussians to densify. (c) The points selected for densification are determined by the substage $k$, $l_{i}$, and viewspace gradients of Gaussians. Residual split is performed upon selected Gaussians, and $l_{j}$ would be assigned to new Gaussians.
3.1 Preliminaries
3D Gaussian Splatting.
The 3D-GS work [23] represents a 3D scene with a set of anisotropic 3D Gaussians for fast and high-quality rendering. Each Gaussian primitive consists of two parameters, mean $\boldsymbol{\mu}$ and covariance matrix $\mathbf{\Sigma}$, representing its position and shape:
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})}.$ (1)
To model the positive semi-definite property of the covariance matrix $\mathbf{\Sigma}$, a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ is used as:
$\mathbf{\Sigma}=\mathbf{RS}\mathbf{S}^{T}\mathbf{R}^{T},$ (2)
where a 3D vector $\mathbf{s}=({s}_{x},{s}_{y},{s}_{y})$ is used for the scaling matrix $\mathbf{S}$.
Furthermore, each Gaussian also contains a spherical harmonics coefficients $SH$ and opacity $o$. In rendering, tile-based rasterization is applied: 3D Gaussians $G(\mathbf{x})$ are projected as 2D Gaussians $G^{\prime}(\mathbf{x})$ on the image plane. A view-dependent color $c$ is then derived from $SH$, and the 2D Gaussians are sorted by depth and rendered using alpha-blending:
$C(\mathbf{x})=\sum_{i\in N_{\mathbf{x}}}c_{i}\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$ (3)
$\alpha_{i}(\mathbf{x})=c_{i}G_{i}^{\prime}(\mathbf{x}),$
where $\mathbf{x}$ is a pixel position in the rendered image, $C(\mathbf{x})$ the color of the pixel, and $N_{\mathbf{x}}$ the number of Gaussians that covers $\mathbf{x}$.
During the training process, 3D Gaussians are periodically densified. 3D-GS calculates the average viewspace gradients $\Delta L_{i}$ over a certain number (e.g. 100) of iterations to determine whether to perform densification. If $\Delta L_{i}$ exceeds threshold $\tau$, a split or clone action is triggered. The split operation addresses over-reconstruction, where excessive space is covered, while clone addresses under-reconstruction, where additional spatial coverage is needed. To determine whether to split or clone, 3D-GS [23] sets a threshold $\tau_{s}=kR$, where $k$ is a predefined factor and $R$ is the scene extent. Split occurs when the scale $s=max({s}_{x},{s}_{y},{s}_{y})$ of a Gaussian surpasses the threshold $\tau_{s}$. Clone is applied when $s<\tau_{s}$.
3.2 ResGS
We propose ResGS, a coarse-to-fine pipeline to boost the rendering quality of 3D-GS with a novel residual densification operation. Fig. 3 shows the overview of our method. We propose a novel densification operation, residual split, (Sec. 3.2.1), to tackle the limitations of the 3D-GS densification operation. Furthermore, we introduce a split-stage supervision schedule to add information to the scene gradually (Sec. 3.2.2). We also design a Gaussian selection scheme that progressively encourages densification for coarse Gaussians to enhance details, as described in Sec. 3.2.3.
3.2.1 Residual Split
The core idea of our densification operation is to generate downscaled Gaussians as residuals to supplement those requiring densification. The newly added Gaussian adaptively enhances the reconstructed model, allowing the scene to be refined as needed. The method is named as residual split. As illustrated in Fig. 3 (a), for a to-be-densified Gaussian $G_{i}$, our densification operation consists of two steps. Firstly, we add a down-scaled replicate $G_{j}$. Specifically, its opacity $o_{j}$, rotation $\mathbf{R}_{j}$, spherical harmonics coefficients $SH_{j}$ would be same as $G_{i}$, $\boldsymbol{\mu}_{j}$ is sampled using the original 3D Gaussian as a PDF for sampling, and $\mathbf{S}_{j}$ is defined as $\mathbf{S}_{i}$ divided by a predefined factor $\lambda_{s}$:
$\mathbf{S}_{j}=\frac{1}{\lambda_{s}}\mathbf{S}_{i},$ (4)
$\mathbf{R}_{j}=\mathbf{R}_{i},\ SH_{j}=SH_{i},\ o_{j}=o_{i},$
$\boldsymbol{\mu}_{j}\sim\mathcal{N}(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}),$
where $\boldsymbol{\Sigma}_{i}$ is the covariance matrix of $G_{i}$. Then, to tackle the density increase in the overlap of the two Gaussians, we reduce the opacity of the original Gaussian. This is accomplished by multiplying its opacity by a predefined factor $\beta$:
$o^{\prime}_{i}=\beta o_{i}.$ (5)
Figure 4: Intermediate results after applying residual split. We saved the Gaussians at 3,000 and 6,000 iterations after applying residual split. As observed, residual split can adaptively handle over-reconstruction and under-reconstruction by refining blurry regions with finer Gaussians while also effectively reconstructing missing structures.
While the clone and split in the original 3D-GS pipeline lack adaptiveness, our residual split can fill in missing geometry and recover details adaptively without requiring a trade-off, as shown in Fig. 2 (c). We further illustrate this using intermediate results, as shown in Fig. 4. In areas where details are challenging to recover, down-scaled Gaussians of finer granularity are progressively introduced to refine and reveal intricate details. On the other hand, since densification does not change the scale of the original Gaussian, the total coverage is increased after densification, helping to reconstruct missing structures. Moreover, our method’s approach to generate nonidentical Gaussians of varying scales enables efficient shape fitting. As illustrated by the blue rectangles in Fig. 1, this non-identicalness allows us to fill textureless regions with a few large Gaussians, significantly reducing redundancy. Consequently, our approach delivers higher rendering quality while using fewer Gaussians.
3.2.2 Image Pyramid as Supervision
To decouple coverage and detail optimization, we design a coarse-to-fine supervision schedule to gradually integrate fine high-frequency information into the scene. Specifically, based on the original multi-view training pictures, we build an image pyramid $\{\mathcal{I}_{i}\}_{i=1}^{L}$ consisting of $L$ layers. $\mathcal{I}_{i}$ corresponds to the set of images at different viewpoints in layer $i$. $\mathcal{I}_{L}$ is the set of original pictures:
$\mathcal{I}_{i}=\{P^{v}_{i}\}_{v\in V},$ (6)
where $V$ is the set of viewpoints and $P^{v}_{i}$ refers to the image at viewpoint $v$ in the $i^{\text{th}}$ layer. Let $(H_{L}^{v},W_{L}^{v})$ be the resolution of $P_{L}^{v}$, and $(H_{i}^{v},W_{i}^{v})$ represent the resolution of $P_{i}^{v}$. Then there is:
$(H_{i}^{v},W_{i}^{v})=(H_{L}^{v}/2^{L-i},W_{L}^{v}/2^{L-i}).$ (7)
Correspondingly, our training process is divided into $L$ stages. At the $i$-th stage, the Gaussians are trained under the supervision of $\mathcal{I}_{i}$. At the initial stage with lower image resolution, the 3D Gaussians are optimized to construct an overall structure of the whole scene from coarse and low-frequency features. As the image resolution increases, high-frequency details of the scene gradually emerge, shifting the focus toward adding finer, more intricate features.
3.2.3 Varying Gradient Threshold
To further emphasize the progressive training, we modify the strategy for selecting Gaussians to densify. As the training proceeds, the training focus is desired to shift toward finer details. Thus, we encourage the densification of coarse Gaussians to capture more detail over time. In our residual split, the generated Gaussians are finer than the original ones. Thus, we can track the relative fineness level of each Gaussian $G_{i}$, denoted as $l_{i}$. We assign $l_{i}$ of all initial Gaussians to 0. If the level of the Gaussian being densified is $l_{i}$, we assign the newly added Gaussian a level of $l_{i}+1$. Thus, as $l_{i}$ gets higher, Gaussians tend to be smaller and finer. Note that $l_{i}$ is only utilized during rendering and discarded after training.
A single stage is then divided evenly into $K$ substages, creating $L\times K$ substages in total. The gradually increasing substage can be seen as an indicator of the scene details level. At substage $k$, we promote additional densification for Gaussians with $l_{i}<k$ and prioritize those at lower levels, achieved by applying varying gradient thresholds. Specifically, the threshold $\tau_{k,i}$ for deciding whether densification is needed for a Gaussian $G_{i}$ is set as:
$\tau_{k,i}=\left\{\begin{aligned} &\tau,l_{i}\geq k,\\ &\frac{\tau}{\alpha^{k-l_{i}}},l_{i}<k,\end{aligned}\right$. (8)
where $\tau$ and $\alpha>1$ are predefined values.
4 Experiments and Results
Table 1: Quantitative comparison to previous methods on real-world datasets. Our method achieves SOTA in the Mip-NeRF360 dataset and comparable results with SOTA on the Tanks&Temples and Deep Blending datasets. indicates results generated by retrained models for different rendering resolution.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem.
Plenoxels 23.08 0.626 0.463 2.1GB 21.08 0.719 0.379 2.3GB 23.06 0.795 0.510 2.7GB
Instant-NGP 25.59 0.699 0.331 48MB 21.92 0.745 0.305 48MB 24.96 0.817 0.390 48MB
Mip-NeRF360 27.69 0.792 0.237 8.6MB 22.22 0.759 0.257 8.6MB 29.40 0.901 0.245 8.6MB
Scaffold-GS 27.69 0.812 0.225 176MB 23.96 0.853 0.177 87MB 30.21 0.906 0.254 66MB
Octree-GS 27.52 0.812 0.218 124MB 24.52 0.866 0.153 84MB 30.41 0.913 0.238 93MB
3D-GS 27.21 0.815 0.214 734MB 23.14 0.841 0.183 411MB 29.41 0.903 0.243 676MB
AbsGS 27.49 0.820 0.191 728MB 23.73 0.853 0.162 304MB 29.67 0.902 0.236  [^4] MB
Pixel-GS 27.52 0.822 0.191 1.32GB 23.78 0.853 0.151 1.06GB 28.91 0.892 0.250 1.09GB
FreGS 27.85 0.826 0.209 - 23.96 0.849 0.178 - 29.93 0.904 0.240 -
Mini-Splatting-D 27.51 0.831 0.176 1.11GB 23.23 0.853 0.140 1.01GB 29.88 0.906 0.211 1.09GB
Ours-Small (AbsGS) 27.94 0.830 0.191 342MB 24.33 0.862 0.150 187MB 30.01 0.906 0.234 289MB
Ours (3D-GS) 28.00 0.831 0.187 600MB 24.26 0.865 0.141 354MB 29.93 0.903 0.232 596MB
Ours (AbsGS) 28.00 0.833 0.174 698MB 24.38 0.867 0.132 351MB 29.91 0.902 0.227 586MB
4.1 Experimental Settings
Following previous works [23, 43, 48, 27, 33, 12], evaluation of our model was done on three datasets, including all nine scenes from Mip-NeRF360 [1], two scenes from Tanks&Temples [24], and two scenes from Deep Blending [17]. The test and train splits follow those of the original 3D-GS [23], and the image resolution is also maintained as in 3D-GS. Three commonly used metrics were used to quantify the fidelity of rendered images: PSNR, SSIM [38], and LPIPS [49]. We also report the memory consumption as an indicator of the number of Gaussians.
We trained three versions of our model: two with AbsGS [43], leveraging its enhanced indication of Gaussians for densification—one standard version and one small version with higher gradient thresholds—and a third version using 3D-GS as the baseline. We applied the periodic opacity reduction from AbsGS across all three versions, as it proved to be a more effective alternative to the opacity reset used in 3D-GS. Additionally, we extended this reduction beyond the densification phase, finding it further minimized redundancy without compromising fidelity. Following previous methods, we trained our model for 30k iterations. The densification stops at 12000 iterations. For the version with AbsGS, $\tau=0.00067$, and $0.0016$ for the small version. For the version with 3D-GS, $\tau=0.00028$. We set $L=3$ and $K=3$. Our first stage lasts for 2500 iterations, the second stage for 3500 iterations, and the rest iterations are in the last stage. $\alpha=2^{\frac{1}{3}}$, $\lambda_{s}=1.6$ and $\beta=0.3$. The loss function remains the same as in 3D-GS [23], consisting of a combination of $\mathcal{L}_{1}$ and D-SSIM losses. All experiments were done on an NVIDIA RTX3090 GPU.
4.2 Comparisons
To assess the effectiveness of our method, we compare it with several 3D-GS variants: 3D-GS [23], Scaffold-GS [27], Octree-GS [33], FreGS [48], AbsGS [43], Pixel-GS [50] and Mini-Splatting-D [12]. Note that Scaffold-GS and Octree-GS use a latent representation rather than the original 3D-GS representation. Some radiance-field-based methods are also selected for comparison, including Plenoxels [45], Instant-NGP [30] and Mip-NeRF360 [1]. We find that the resolution of Mip-NeRF360 dataset images varies across different works. Following the resolution setup in the 3D-GS work, we retrain some models on the Mip-NeRF360 dataset. Additionally, for works that did not report their storage usage, we report the results based on their released code.
Quality Comparisons.
In Table 1, we present quantitative results comparing the rendering quality of our method with other approaches. From the table, we can see that our method achieves SOTA performance across all metrics on the Mip-NeRF360 [1] dataset. On the Tanks&Temples [24] dataset, our method also achieves the best SSIM and LPIPS metrics, which closely reflect human perception. In the Deep Blending [17] dataset, though our method does not achieve the highest score, it shows a notable improvement over the baseline. Additionally, our model variant without AbsGS [43] maintains similar performance to the full model, with only a slight reduction in LPIPS. Notably, our smaller model version achieves strong performance with reduced memory usage, showcasing the efficiency and scalability of our approach. In Fig. 5, we present qualitative comparisons between our method and several 3D-GS variants [23, 50, 43] across various scenes, including both indoor and outdoor settings. It is evident that our model captures finer details, as shown in the second, third, and fourth rows, constructs missing geometry more effectively (last row), and achieves more accurate geometry reconstruction (first row).
Figure 5: Qualitative comparisons of ResGS with three 3D-GS variations on a variety of indoor and outdoor scenes. Our approach captures more intrinsic details and acquires more complete geometry in complex scenes.
Table 2: Efficiency comparison with former methods.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
FPS Time FPS Time FPS Time
Octree-GS 119 29m55s 126 23m8s 151 34m38s
Scaffold-GS 103 23m53s 137 15m37s 141 19m40s
3D-GS 103 27m16s 127 15m34s 101 24m48s
AbsGS 132 28m48s 193 13m35s 185 22m3s
Mini-Splatting-D 74 34m42s 75 27m39s 100 30m37s
Pixel-GS 65 40m6s 70 29m29s 70 34m1s
Ours-Small (AbsGS) 141 20m56s 206 11m35s 202 19m26s
Ours (AbsGS) 100 29m24s 150 16m29s 153 23m33s
Table 3: Analyzing residual split upon multiple 3D-GS variants. The performance consistently increases on these pipelines, indicating that our residual split is a better densification operation. 3D-GS indicates a retrained version of 3D-GS [23] for better performance.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem
3D-GS 27.21 0.815 0.214 734MB 23.14 0.841 0.183 411MB 29.41 0.903 0.243 676MB
3D-GS 27.38 0.813 0.215 821MB 23.66 0.844 0.179 434MB 29.48 0.900 0.247 663MB
3D-GS + residual split 27.44 0.813 0.216 586MB 23.76 0.845 0.178 318MB 29.75 0.902 0.240 546MB
AbsGS 27.49 0.820 0.191 728MB 23.73 0.853 0.162 304MB 29.67 0.902 0.236  [^4] MB
AbsGS + residual split 27.71 0.827 0.185 712MB 24.05 0.857 0.161 285MB 29.68 0.905 0.234 421MB
Pixel-GS 27.52 0.822 0.191 1.32GB 23.78 0.853 0.151 1.06GB 28.91 0.892 0.250 1.09GB
Pixel-GS + residual split 27.62 0.821 0.191 1.00GB 24.05 0.856 0.148 754MB 29.75 0.899 0.235 914MB
Mini-Splatting-D 27.51 0.831 0.176 1.11GB 23.23 0.853 0.140 1.01GB 29.88 0.906 0.211 1.09GB
Mini-Splatting-D + residual split 27.64 0.831 0.176 1.04GB 23.49 0.853 0.140 967MB 29.94 0.903 0.210 1.02GB
Efficiency Comparisons.
We provide the computational costs of our method in comparison with other approaches in Table 2. It shows that our method involves a slight increase in training time while delivering much higher rendering quality than methods with lower rendering quality like 3D-GS [23] and AbsGS [43]. Our method is also significantly more efficient than other methods like Mini-Splatting-D [12], with comparable rendering quality. Notably, our small model provides the fastest FPS and shortest training time in all compared methods while maintaining high fidelity.
4.3 Analysis
4.3.1 Residual Split
Compatibility.
We analyze the compatibility of our novel densification operation by replacing split and clone in multiple 3D-GS variants [23, 43, 50, 12] with residual split. The hyperparameters are set the same as in the original pipelines. As Table 3 shows, our method leads to a notable increase in PSNR metrics across all pipelines, while for other metrics, some exhibit improvements, and others remain approximately the same. Additionally, memory consumption is reduced after incorporating residual split compared to the original pipeline. This underscores our method’s ability to improve rendering quality and reduce redundancy, while also demonstrating its compatibility and potential for broader application across 3D-GS-oriented works as an effective alternative to split and clone.
Figure 6: Impact of the opacity reduction weight $\beta$ for residual split.
Figure 7: Impact of the scale reduction factor $\lambda_{s}$ for residual split.
Hyperparameter Sensitivities.
Our residual split involves some hyperparameters: opacity reduction weight $\beta$ and scale reduction factor $\lambda_{s}$. We conduct a corresponding sensitivity analysis on the Tanks&Temples [24] dataset while applying different $\beta\in[0.05,1.0)$ and $\lambda_{s}\in[1.05,2.0]$. All models are trained with similar storage. The LPIPS scores with different parameters are shown in Fig. 6 and 7. As they show, the LPIPS score remains relatively stable when $\beta<0.4$ and $\lambda_{s}>1.55$, indicating that our method is robust to hyperparameter variations, with only a 0.006 difference in LPIPS between the best and worst cases and a relatively broad range where rendering quality remains optimal.
4.3.2 Ablation Studies
Table 4: Ablation Study on our pipeline. IP denotes image pyramid supervision, RS represents using residual split for densification, and VT indicates applying varying gradient threshold. Each component contributes to our method’s effectiveness.
Dataset Mip-NeRF360 Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Base 27.41 0.817 0.189 29.60 0.899 0.240
Base + RS 27.66 0.825 0.183 29.68 0.900 0.233
Base + IP 27.54 0.823 0.182 29.05 0.896 0.245
Base + RS + IP 27.88 0.831 0.178 29.82 0.902 0.230
Base + RS + IP + VT (full) 28.00 0.833 0.174 29.91 0.902 0.227
To assess the effectiveness of our pipeline’s key modules, including residual split, image pyramid as supervision, and the varying gradient threshold, we conduct ablation studies on the Mip-NeRF360 [1] and Deep Blending [17] datasets. For the varying gradient threshold, we only conduct studies on its removal, as it is highly interdependent with the other two modules. As the results in Table 4 shows, both residual split and image pyramid supervision individually boost the baseline’s performance on the Mip-NeRF360 dataset.
However, on Deep Blending, we observe a performance drop in Base+IP. This is because the dataset contains many weakly textured areas. As noted in Sec. 1, split and clone tend to fill such regions with an excessive number of small Gaussians. This makes them prone to overfitting the training views and lead to reduced performance on test views. The issue is amplified in the early iterations of a coarse-to-fine schedule, where only coarse scene information is available, making weakly textured areas even harder to resolve. As a result, these small Gaussians struggle to construct accurate geometry, leading to the performance drop observed in Base+IP.
Conversely, when using RS, weakly textured areas in early iterations are represented by only a few Gaussians, reducing overfitting. In later iterations, as finer information becomes available, RS progressively introduces more small Gaussians. As a result, RS benefits from progressive training, as evidenced by the improved performance of Base+RS+IP over Base+RS. This further proves RS is more effective than split and clone, especially in a coarse-to-fine schedule. Additionally, applying the varying gradient threshold (our full model) further boosts performance on both datasets, validating its effectiveness.
4.4 Discussions and Limitations
The above results demonstrate that our method significantly enhances rendering quality on the Mip-NeRF360 [1] and Tanks&Temples [24] datasets, achieving state-of-the-art performance. However, for the Deep Blending dataset [17], our method fails to achieve the best performance on every metric. While our approach implicitly constrains the number of Gaussians in weakly textured regions, which helps mitigate overfitting and improves quality, it does not include explicit regularization. This suggests an opportunity for further optimization in future work.
5 Conclusion
In this paper, we analyze the limitations of the current 3D-GS densification operation. To address these issues, we propose residual split, a residual empowered densification operation that adaptively handles both over-reconstruction and under-reconstruction. Additionally, we introduce ResGS, a coarse-to-fine 3D-GS pipeline to supplement residual split. Specifically, we use an image pyramid for supervision to gradually introduce fine details during Gaussian optimization. We also design a Gaussian selection scheme to further densify coarse Gaussians progressively. Our pipeline achieves SOTA rendering quality on several datasets while our residual split is highly compatible with other 3D-GS variants.
Acknowledgements
This work was supported by CAS Project for Young Scientists in Basic Research (YSBR-116).
Supplementary Material
6 Additional Experiment Results
In this section, we provide additional experimental results of our work.
6.1 Evaluation on large-scale urban datasets
To evaluate the performance of our approach on large-scale urban datasets, we performed experiments using the Waymo [36] dataset and compared our results against GaussianPro [11]. The results are shown in the table below. It can be seen that our method achieves better results across all metrics.
Table 5: Results on the Waymo dataset compared with GaussianPro.
Dataset Waymo
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem.
GaussianPro 34.68 0.949 0.191 297MB
Ours 35.29 0.951 0.191 284MB
6.2 Variance of PSNR on Deep Blending Dataset
A notable variance in the PSNR metrics of our method is observed on the Deep Blending [17] dataset. We tested our method on the Deep Blending dataset 10 times and reported the average performance and population standard deviation. The results are shown in Table 6. The table demonstrates that the PSNR scores on the Deep Blending dataset exhibit a notable variance. The underlying reason, as discussed in our limitations, is that our method does not address issues related to occlusion and overfitting, making the Deep Blending dataset particularly challenging for our approach. This highlights opportunities for future optimization and improvement. Furthermore, we report the average performance and population standard deviation over 10 runs for the other two datasets, Mip-NeRF360 [1] and Tanks&Temples [24], as presented in Table 7 and Table 8, respectively. The PSNR metrics for these two datasets do not show a notable variance, showcasing the robustness of our method in most scenarios.
Table 6: Average and standard deviation results of three metrics on the Deep Blending dataset.
Dataset Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Avg Stdev Avg Stdev Avg Stdev
Ours-Small (AbsGS) 29.71 0.211 0.903 0.002 0.235 0.002
Ours (3D-GS) 29.64 0.205 0.900 0.003 0.233 0.002
Ours (AbsGS) 29.68 0.173 0.900 0.002 0.228 0.002
Table 7: Average and standard deviation results of three metrics on the Mip-NeRF360 dataset.
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Avg Stdev Avg Stdev Avg Stdev
Ours-Small (AbsGS) 27.93 0.010 0.830 0.001 0.191 0.001
Ours (3D-GS) 27.99 0.011 0.831 0.001 0.187 0.001
Ours (AbsGS) 27.99 0.011 0.833 0.001 0.174 0.001
Table 8: Average and standard deviation results of three metrics on the Tank&Temples dataset.
Dataset Tank&Temples
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Avg Stdev Avg Stdev Avg Stdev
Ours-Small (AbsGS) 24.21 0.081 0.862 0.001 0.151 0.001
Ours (3D-GS) 24.26 0.058 0.864 0.001 0.141 0.001
Ours (AbsGS) 24.27 0.063 0.866 0.001 0.133 0.001
6.3 Further Analysis of Residual Split on 3D-GS
Table 9: Results of residual split upon 3D-GS with varied gradient threshold.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem
3D-GS 27.21 0.815 0.214 734MB 23.14 0.841 0.183 411MB 29.41 0.903 0.243 676MB
3D-GS + residual split 27.48 0.815 0.213 722MB 23.85 0.846 0.174 374MB 29.83 0.901 0.238 640MB
In our paper, we tested the compatibility of our residual split across multiple pipelines without modifying any hyper-parameters. Here, we present the results of applying residual split to 3D-GS [23] while adjusting the densification threshold $\tau$ to maintain the same storage size as 3D-GS, further demonstrating the capability of our method to enhance rendering quality. The results are shown in Table 9, showing that our method yields a further performance increase.
Table 10: Results of our method without extended periodic opacity reduction. Here, No EPOR refers to our method trained without applying the extended periodic opacity reduction. No EPOR + CT refers to the model without the extended periodic opacity reduction but with an adjusted gradient threshold to ensure that the final storage size is approximately the same as our full model with extended periodic opacity reduction.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem
Ours (No EPOR) 28.02 0.833 0.173 973MB 24.32 0.867 0.130 591MB 29.86 0.902 0.227 911MB
Ours (No EPOR + CT) 28.00 0.833 0.181 608MB 24.35 0.866 0.139 380MB 29.96 0.904 0.230 564MB
Ours (full) 28.00 0.833 0.174 698MB 24.38 0.867 0.132 351MB 29.91 0.902 0.227 586MB
6.4 Results using random initialization
To assess the robustness of our method to random initialization, we conducted experiments using randomly initialized point clouds and compared our approach with other methods trained on the same random point clouds. The results are shown in Table 11. For the Mip-NeRF360 [1] and Tanks&Temples [24] datasets, our model shows a noticeable drop in quality compared to when initialized with SfM [35]. However, compared to other methods, it still achieves the best performance across all three metrics on Mip-NeRF360, and the highest SSIM and LPIPS on Tanks&Temples. As for Deep Blending [17], our method and others, excluding Scaffold-GS and Octree-GS, do not exhibit a significant performance drop. While we do not achieve the highest scores, our method remains competitive, ranking second across all metrics.
Table 11: Results using random initialized point clouds, compared with other methods.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem.
Scaffold-GS 25.90 0.743 0.301 157MB 22.00 0.768 0.282 51MB 29.48 0.891 0.285 56MB
Octree-GS 25.70 0.739 0.292 190MB 22.15 0.794 0.226 81MB 28.29 0.873 0.300 92MB
3D-GS 25.90 0.769 0.266 632MB 21.19 0.777 0.241 339MB 29.42 0.897 0.253 605MB
AbsGS 25.50 0.773 0.226 741MB 20.89 0.786 0.211 304MB 29.52 0.897 0.241 478MB
Mini-Splatting-D 25.18 0.792 0.207 1.10GB 19.77 0.760 0.220 1.02GB 29.76 0.901 0.212 1.09GB
Pixel-GS 26.23 0.792 0.221 1.14GB 21.10 0.788 0.207 939MB 28.76 0.889 0.256 1.04GB
Ours-Random 26.38 0.799 0.200 750MB 22.02 0.815 0.178 282MB 29.57 0.899 0.234 464MB
Ours-SfM 28.00 0.833 0.174 698MB 24.38 0.867 0.132 351MB 29.91 0.902 0.227 586MB
6.5 Periodic Opacity Reduction
3D-GS [23] often encounters redundancy issues, which we categorize into two main causes: the first is the use of an excessive number of small-scale Gaussians to represent coarse areas, and the second is the overlap between Gaussians, resulting in redundant Gaussians that contribute minimally to the rendered image. Our residual split can solve the first issue but does not address the second. To tackle the second issue, as mentioned in our experiment settings, we implemented the periodic opacity reduction technique in [43, 5]. Specifically, the opacity of Gaussians is periodically reduced, causing the opacity of those that contribute little to the rendered image to diminish to a negligible value, allowing them to be effectively pruned. Additionally, we extended this operation to occur after the densification stage, discovering that it further reduces redundancy without compromising fidelity. Here, we present the experimental results of our method without extending the periodic opacity reduction operation, as shown in Table 10. From the table, we can observe that extending the periodic opacity reduction operation effectively reduces redundancy without compromising the fidelity of the rendered images. Furthermore, when the operation is not extended and the gradient threshold is adjusted to maintain storage approximately equal to our full model, the rendering quality shows only a slight decrease in LPIPS metrics. This indicates that while extending the operation can enhance the performance of our method, its overall impact on the performance improvement is relatively minor.
6.6 Image Resolution of Mip-NeRF360 Dataset
In our paper, we mentioned that the image resolution on the Mip-NeRF360 [1] dataset varies across different works. 3D-GS [23] uses the provided “images_4”, the official downscaled 4 times images from the dataset, for outdoor scenes and “images_2” for indoor scenes. Most works [43, 12, 48] follow the 3D-GS setting, but Scaffold-GS [27], Octree-GS [33], Pixel-GS [50] uses different settings. Specifically, Scaffold-GS and Octree-GS downscale original images to 1.6K resolution, while Pixel-GS downscales original images by a factor of 4 for outdoor scenes and 2 for indoor scenes during the training process, instead of using the provided downscaled images. This discrepancy results in differences in experimental settings. To further demonstrate the performance of our method, we provide its results under the two different settings, as shown in Table 12 and 13. The tables show that our method prevails even under these alternate settings.
Table 12: Results on the Mip-NeRF360 dataset with same setting as Scaffold-GS and Octree-GS.
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Scaffold-GS 27.71 0.813 0.221
Octree-GS 27.73 0.815 0.217
Ours-Small (AbsGS) 28.04 0.831 0.191
Ours (3D-GS) 28.10 0.832 0.187
Ours (AbsGS) 28.14 0.834 0.174
Table 13: Results on the Mip-NeRF360 dataset with same setting as Pixel-GS.
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Pixel-GS 27.88 0.834 0.176
Ours-Small (AbsGS) 28.21 0.841 0.176
Ours (3D-GS) 28.28 0.843 0.172
Ours (AbsGS) 28.30 0.845 0.160
6.7 Per Scene Results
We present the per-scene results of the used metrics in Table 14-19. For works that did not report per-scene results, we obtained and reported them using their released code. The works selected for comparison are mainly the same as in the paper: 3D-GS [23], Scaffold-GS [27], Octree-GS [33], AbsGS [43], Pixel-GS [50], Mini-Splatting-D [12], Plenoxels [45], Instant-NGP [30] and Mip-NeRF360 [1]. Note that we did not show the results of FreGS [48] since they did not report their per-scene results or release their code. Works retrained on the Mip-NeRF360 [1] are marked with a .
Table 14: Per-scene PSNR metrics of Tank&Temples and Deep Blending dataset.
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 23.22 18.93 23.14 22.98
Instant-NGP 23.38 20.46 28.26 21.67
Mip-NeRF360 24.91 19.52 29.14 29.66
Scaffold-GS 25.77 22.15 29.80 30.62
Octree-GS 26.27 22.77 29.87 30.95
3D-GS 25.19 21.10 28.77 30.04
AbsGS 25.74 21.72 29.20 30.14
Pixel-GS 25.49 22.13 28.02 29.79
Mini-Splatting-D 25.43 21.04 29.32 30.43
Ours-Small (AbsGS) 26.07 22.40 29.64 30.38
Ours (3D-GS) 26.08 22.65 29.54 30.32
Ours (AbsGS) 26.14 22.61 29.51 30.32
Table 15: Per-scene SSIM metrics of Tank&Temples and Deep Blending dataset.
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 0.774 0.663 0.787 0.802
Instant-NGP 0.800 0.689 0.854 0.779
Mip-NeRF360 0.857 0.660 0.901 0.900
Scaffold-GS 0.883 0.822 0.907 0.904
Octree-GS 0.896 0.835 0.912 0.914
3D-GS 0.879 0.802 0.899 0.906
AbsGS 0.888 0.818 0.898 0.907
Pixel-GS 0.883 0.823 0.885 0.899
Mini-Splatting-D 0.890 0.817 0.905 0.908
Ours-Small (AbsGS) 0.889 0.834 0.905 0.907
Ours (3D-GS) 0.890 0.840 0.901 0.905
Ours (AbsGS) 0.893 0.841 0.900 0.905
Table 16: Per-scene LPIPS metrics of Tank&Temples and Deep Blending dataset.
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 0.335 0.422 0.521 0.499
Instant-NGP 0.249 0.360 0.352 0.428
Mip-NeRF360 0.159 0.354 0.237 0.252
Scaffold-GS 0.147 0.206 0.250 0.258
Octree-GS 0.118 0.198 0.231 0.244
3D-GS 0.148 0.218 0.244 0.241
AbsGS 0.132 0.193 0.241 0.233
Pixel-GS 0.122 0.180 0.258 0.243
Mini-Splatting-D 0.100 0.181 0.218 0.204
Ours-Small (AbsGS) 0.124 0.177 0.233 0.235
Ours (3D-GS) 0.118 0.164 0.232 0.231
Ours (AbsGS) 0.106 0.159 0.231 0.223
Table 17: Per-scene PSNR metrics of Mip-NeRF360 dataset.
Method Scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Plenoxels 21.91 20.10 23.49 20.66 22.25 27.59 23.62 23.42 24.67
Instant-NGP 22.17 20.65 25.07 23.47 22.37 29.69 26.69 29.48 30.69
Mip-NeRF360 24.31 21.65 26.88 26.18 22.93 31.47 29.45 31.99 33.40
Scaffold-GS 25.07 21.36 27.35 26.64 22.93 32.04 29.50 31.63 32.71
Octree-GS 25.05 21.33 27.58 26.41 22.81 32.09 29.48 30.89 31.71
3D-GS 25.25 21.52 27.41 26.55 22.49 30.63 28.70 30.32 31.98
AbsGS 25.29 21.35 27.48 26.71 21.99 31.61 29.03 31.62 32.32
Pixel-GS 25.21 21.49 27.42 26.84 22.09 31.45 29.05 31.65 32.53
Mini-Splatting-D 25.55 21.50 27.67 27.11 22.13 31.41 28.72 31.75 31.72
Ours-Small (AbsGS) 25.62 21.80 27.72 27.27 22.90 32.41 29.32 31.95 32.48
Ours (3D-GS) 25.55 22.07 27.71 27.12 22.82 32.48 29.41 32.09 32.74
Ours (AbsGS) 25.62 21.78 27.82 27.19 22.57 32.51 29.50 32.26 32.78
Table 18: Per-scene SSIM metrics of Mip-NeRF360 dataset.
Method Scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Plenoxels 0.496 0.431 0.6063 0.523 0.509 0.8417 0.759 0.648 0.814
Instant-NGP 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906
Mip-NeRF360 0.685 0.584 0.809 0.745 0.631 0.910 0.892 0.917 0.938
Scaffold-GS 0.755 0.590 0.858 0.765 0.639 0.923 0.910 0.926 0.943
Octree-GS 0.759 0.597 0.864 0.763 0.643 0.924 0.907 0.916 0.930
3D-GS 0.771 0.605 0.868 0.775 0.638 0.914 0.905 0.922 0.938
AbsGS 0.783 0.623 0.871 0.780 0.617 0.925 0.911 0.929 0.945
Pixel-GS 0.775 0.633 0.867 0.784 0.630 0.920 0.911 0.929 0.944
Mini-Splatting-D 0.798 0.642 0.878 0.804 0.640 0.928 0.913 0.934 0.946
Ours-Small (AbsGS) 0.792 0.634 0.872 0.802 0.652 0.929 0.914 0.932 0.945
Ours (3D-GS) 0.790 0.639 0.874 0.800 0.650 0.929 0.916 0.934 0.947
Ours (AbsGS) 0.797 0.647 0.876 0.804 0.645 0.931 0.918 0.934 0.948
Table 19: Per-scene LPIPS metrics of Mip-NeRF360 dataset.
Method Scenes
bicycle flowers garden stump treehill room counter kitchen bonsai
Plenoxels 0.506 0.521 0.3864 0.503 0.540 0.4186 0.441 0.447 0.398
Instant-NGP 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205
Mip-NeRF360 0.301 0.344 0.170 0.261 0.339 0.211 0.204 0.127 0.176
Scaffold-GS 0.233 0.352 0.120 0.236 0.331 0.216 0.203 0.130 0.207
Octree-GS 0.225 0.340 0.108 0.233 0.310 0.203 0.202 0.141 0.219
3D-GS 0.205 0.336 0.103 0.210 0.317 0.220 0.204 0.129 0.205
AbsGS 0.171 0.270 0.100 0.195 0.278 0.200 0.189 0.121 0.190
Pixel-GS 0.182 0.263 0.100 0.186 0.280 0.213 0.185 0.120 0.193
Mini-Splatting-D 0.158 0.255 0.090 0.169 0.262 0.190 0.172 0.114 0.175
Ours-Small (AbsGS) 0.175 0.290 0.099 0.188 0.279 0.196 0.184 0.117 0.189
Ours (3D-GS) 0.168 0.295 0.092 0.179 0.280 0.192 0.177 0.114 0.185
Ours (AbsGS) 0.156 0.251 0.089 0.170 0.250 0.187 0.172 0.112 0.179